# The possible adelic indices for elliptic curves admitting a rational cyclic isogeny

### v1

@@ 8.3. Rational Isogenies of Degree 33
Let $E/\mathbb{Q}$ be a non-CM elliptic curve that admits a rational isogeny of degree $3$ and no rational isogeny of any larger prime degree. By Proposition 20, we know all possible $2$-adic and $3$-adic Galois images of $E$ that contain I-I; we denote the sets of these images by $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$, respectively. In particular, $\mathcal{L}_{2}$ consists of $15$ groups and $\mathcal{L}_{3}$ consists of $14$ groups. There are two possibilities for the $5$-adic image: $\operatorname{GL}_{2}(\mathbb{Z}_{5})$ and 5.5.0.a.1. However, we note that the modular curve 15.20.1.a.1 $X_{0}(3)\times$ 5.5.0.a.1 is an elliptic curve of rank 0. It has only two noncuspidal non-CM rational points. These give rise only to elliptic curves of adelic index $120$. For this reason, we will assume the $5$-adic Galois representation is surjective for the remainder of the analysis. Then, by Proposition 20, the $\ell$-adic Galois representation of $E$ is surjective for all $\ell\geq 5$.
@@
We now discuss the computation of their rational points. All of the computational claims about rational points for this section can be verified in the file Section 8.3.m.
Among the $40$ curves, the following are elliptic curves of rank 0: 6.24.1.a.1, 12.24.1.b.1, 12.72.1.d.1, 12.72.1.f.1, 18.24.1.a.1, 72.24.1.a.1, 72.72.1.d.1, 24.24.1.bx.1, 24.72.1.t.1, 36.24.1.a.1, 36.72.1.b.1, and 36.72.1.c.1. Their only rational points are either cusps or CM points. The elliptic curve 12.32.1.b.1 is of rank 0 and has eight rational points. Its non-CM $j$-invariants are $\{\tfrac{-3^{3}\cdot 11^{3}}{2^{2}},\tfrac{3^{2}\cdot 23^{3}}{2^{6}}\}$. The adelic index for both of these $j$-invariants is $128\in\mathcal{I}_{3}$.
The following genus 2 curves are all rank 0: 54.108.2.a.1, 72.72.2.a.1, 18.36.2.d.1, 18.72.2.b.1, 18.72.2.b.2, 18.72.2.d.1, 18.108.2.a.1, 18.108.2.b.1, 72.72.2.a.2, 72.72.2.b.1, 72.72.2.b.2, 72.72.2.c.1, 18.108.2.c.1, 18.108.2.d.1, 18.108.2.d.2, 18.108.2.e.1, 36.72.2.a.1,
36.72.2.a.2, 36.72.2.b.1. Using Chabauty0, we compute all the rational points. All of their rational points, if any exist, are either cusps or CM points.
Next, let $C$ denote the subscheme of $\mathbb{P}^{3}_{\mathbb{Q}}$ defined by the sequence of homogeneous polynomials that represent the canonical model of the genus 2 curve 72.72.2.d.1. We verify that the set of points $C(\mathbb{Z}/4\mathbb{Z})$ is empty. Thus, there are no rational points.
@@
There are a total of $17$ curves for which a fiber product could possibly give rise to an adelic index of $432$ or $3888$. These are:
@@
We now discuss their rational points. The curve 6.24.1.b.1 is an elliptic curve of rank 0. It has exactly 6 rational points, all of which are cusps or CM points. The curve 18.24.1.b.1 is also an elliptic curve of rank 0, and has 6 points all of which are cusps or CM points. The curve 18.72.1.e.1 has 3 points, all of which are cusps or CM points. The curve 54.72.3.d.1, which we denote by $C$, has genus 3, rank 1. It has a canonical model in $\mathbb{P}^{2}$ given by
@@
We will use the argument of Section 3.2. The curve $C$ has at least three rational points. Set $P_{1}:=(1,0,0)$, $P_{2}:=(0,1,0)$ and $P_{3}:=(0,0,1)$. It has an automorphism $x,y,z]\mapsto[x,z,y]$. The quotient is a rank 1 elliptic curve. Therefore, there exists an abelian variety of rank 0, which we denote $V$, such that the Jacobian of this curve decomposes as $E\times V$. For any rational point $P$ on this curve, the point Pi(P)P-i(P) lies in the torsion subgroup of its Jacobian, which we denote by $J_{C}(\mathbb{Q})_{\operatorname{tors}}$. It suffices to compute preimages of rational points in $J_{C}(\mathbb{Q})_{\operatorname{tors}}$ under the map $a\colon C(\mathbb{Q})\to J_{C}(\mathbb{Q})_{\operatorname{tors}}$ given by $P\mapsto P-i(P)$. Note that $a$ is injective away from the fixed points of $i$. By using local computations, we can deduce that $J_{C}(\mathbb{Q})_{\operatorname{tors}}$ is a subgroup of $\mathbb{Z}/3\mathbb{Z}\times\mathbb{Z}/3\mathbb{Z}$. Further, we observe that the divisors $D_{1}:=P_{1}-P_{2}$ and $D_{2}:=P_{2}-P_{3}$ are of order 3 and generate $J_{C}(\mathbb{Q})_{\operatorname{tors}}$. One can see that $P_{2}-i(P_{2})$ is equivalent to $D_{2}$ and $P_{3}-i(P_{3})$ is equivalent to $2D_{2}.$ Further, by working modulo $5$, we check that there is no divisor of the form Pi(P)P-i(P) which is equivalent to $aD_{1}+bD_{2}$ for $(a,b)$ other than $(0,1)$ and $(0,2).$ Therefore, the only rational points of $C$ are the fixed points of $i$ and $\{P_{2},P_{3}\}$. It is easy to check that the only fixed point is $P_{1}$. This means that the only rational points are the known ones. Two are cusps and one is a CM point. The curves 54.72.4.e.1 and 54.72.5.d.1 have a map to some rank 0 elliptic curve. Using the code ModCrvToEC [JacobJeremycode] we determine that they have four and three points respectively. All are either cusps or CM. The curves 18.72.3.j.1, 18.72.3.n.1 and 18.72.3.m.1 cover 6.24.1.b.1, which only has cusps and CM points. The curves 18.72.3.k.1, 18.72.3.k.2, 18.72.3.o.1 and 54.72.3.c.1 cover 18.24.1.b.1, which has exactly six points, all of which are cusps or CM points. The curves 18.72.2.e.1 and 18.72.2.e.2 are genus 2 rank 0. Using Chabauty0, we find that they both have 3 points, and all are cusps. The curve 18.72.2.f.1 has genus 2, rank 2, and is bielliptic with Weierstrass model

### v2

@@ 8.3. Rational Isogenies of Degree 33
Let $E/\mathbb{Q}$ be a non-CM elliptic curve that admits a rational isogeny of degree $3$ and no rational isogeny of any larger prime degree. By Proposition 19, we know all possible $2$-adic and $3$-adic Galois images of $E$ that contain I-I; we denote the sets of these images by $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$, respectively. In particular, $\mathcal{L}_{2}$ consists of $15$ groups and $\mathcal{L}_{3}$ consists of $14$ groups. There are two possibilities for the $5$-adic image: $\operatorname{GL}_{2}(\mathbb{Z}_{5})$ and 5.5.0.a.1. However, we note that the modular curve 15.20.1.a.1 $X_{0}(3)\times$ 5.5.0.a.1 is an elliptic curve of rank 0. It has only two noncuspidal non-CM rational points. These give rise only to elliptic curves of adelic index $160$. For this reason, we will assume the $5$-adic Galois representation is surjective for the remainder of the analysis. Then, by Proposition 19, the $\ell$-adic Galois representation of $E$ is surjective for all $\ell\geq 5$.
8.3.1. Maximal Products with Finitely Many Rational Points
@@
As before, we will only highlight unusual cases in this article. Full computations can be found on the GitHub repository. In particular, all of the computational claims about rational points for this section can be verified in the file Section 8.3.m.
The elliptic curve 12.32.1.b.1 is of rank 0 and has eight rational points. Its non-CM $j$-invariants are $\{\tfrac{-3^{3}\cdot 11^{3}}{2^{2}},\tfrac{3^{2}\cdot 23^{3}}{2^{6}}\}$. The adelic index for both of these $j$-invariants is $128\in\mathcal{I}_{3}$.
Let $C$ denote the subscheme of $\mathbb{P}^{3}_{\mathbb{Q}}$ defined by the sequence of homogeneous polynomials that represent the model of the genus 2 curve 72.72.2.d.1 produced by Zywina’s code. We verify that the set of points $C(\mathbb{Z}/4\mathbb{Z})$ is empty. Thus, there are no rational points.
@@
8.3.2. Products with Infinitely many Rational Points
@@
There are a total of $17$ curves for which a fiber product could possibly give rise to an adelic index of $432$ or $3888$. Of these, for the following $15$ curves we need to rule out the adelic index of $432$:
For the remaining two curves, 9.36.0.b.1 and 27.36.0.a.1, we need to rule out the adelic index of $3888$.
@@
These are mostly also amenable to techniques already discussed in this paper. We go through each in full detail in Section 8.3.m. We provide details about the curves with ad hoc approaches here.
The curve 54.72.3.d.1, which we denote by $C$, has genus 3, rank 1. It has a canonical model in $\mathbb{P}^{2}$ given by
@@
We will use the argument of Section 3.2. The curve $C$ has at least three rational points. Set $P_{1}:=(1:0:0)$, $P_{2}:=(0:1:0)$ and $P_{3}:=(0:0:1)$. It has an automorphism $x:y:z]\mapsto[x:z:y]$. The quotient is a rank 1 elliptic curve. Therefore, there exists an abelian variety of rank 0, which we denote $V$, such that the Jacobian of this curve decomposes as $E\times V$. For any rational point $P$ on this curve, the point Pi(P)P-i(P) lies in the torsion subgroup of its Jacobian, which we denote by $J_{C}(\mathbb{Q})_{\operatorname{tors}}$. It suffices to compute preimages of rational points in $J_{C}(\mathbb{Q})_{\operatorname{tors}}$ under the map $a\colon C(\mathbb{Q})\to J_{C}(\mathbb{Q})_{\operatorname{tors}}$ given by $P\mapsto P-i(P)$. Note that $a$ is injective away from the fixed points of $i$. By using local computations, we can deduce that $J_{C}(\mathbb{Q})_{\operatorname{tors}}$ is a subgroup of $\mathbb{Z}/3\mathbb{Z}\times\mathbb{Z}/3\mathbb{Z}$. Further, we observe that the classes of the divisors $D_{1}:=P_{1}-P_{2}$ and $D_{2}:=P_{2}-P_{3}$ are of order 3 and generate $J_{C}(\mathbb{Q})_{\operatorname{tors}}$. One can see that $P_{2}-i(P_{2})$ is equivalent to $D_{2}$ and $P_{3}-i(P_{3})$ is equivalent to $2D_{2}.$ Further, by working modulo $5$, we check that there is no divisor of the form Pi(P)P-i(P) which is equivalent to $aD_{1}+bD_{2}$ for $(a,b)$ other than $(0,1)$ and $(0,2).$ Therefore, the only rational points of $C$ are the fixed points of $i$ and $\{P_{2},P_{3}\}$. It is easy to check that the only fixed point is $P_{1}$. This means that the only rational points are the known ones. Two are cusps and one is a CM point.
The curves 54.72.4.e.1 and 54.72.5.d.1 have a map to some rank 0 elliptic curve. Using the code ModCrvToEC [31] we determine that they have four and three rational points, respectively. All are either cusps or CM.
The curve 18.72.2.f.1 has genus 2, rank 2, and is bielliptic with Weierstrass model